In the laboratory a coffee cup calorimeter, or constant pressure calorimeter, is frequently used to determine the specific
Answer #1

Heat capacity of water = 4.184 J/(g. °C) Heat capacity of Cu = 0.385 J/(g. °C) qwater = m* C* AT = 80.89 g * 4.184 J/(g. °C)